steve travels 4 times as fast as fred. traveling in opposite directions, they are 105 miles apart after 3 hours. find their rates of travel.
Answer by stanbon(75887)

steve travels 4 times as fast as fred. traveling in opposite directions, they are 105 miles apart after 3 hours. find their rates of travel.
=========
Steve DATA"
rate 4x ; time = 3 hrs ; distance = 3*4x = 12x miles
------
Fred DATA:
rate = x ; time = 3 hrs ; distance 3x miles
-------
Equation:
dist + dist = 105 miles
12x + 3x = 105
-----
15x = 105
x = 7 mph (Fred's rate)
4x = 28 mph (Steve's rate)
---------
Cheers,
Stan H.
-----------